NDOS Commissioning Meeting 2011/04/18

Attendance: Jon, Rafael, Kurt, Leon, Athans, Andrew, Rick K., Enrique, Sue, Jarek, Minerba, Mark Mathis, Tricia


  1. NDOS Status / Installation Plans
  2. DAQ/Electronics Status & Plans
  3. DCS Status & Plans
  4. AOB

NDOS Status / Installation Plans

  • Leon: cap-boards should be arriving this week (probably on-site already), will want to install those on 1st 2 di-blocks.
  • Andrew:
    • change control went thru on networking, we are getting new switch in LCC today, will go live tomorrow. More ports, more nodes.
    • suggestion is that 1 switch is used for controls only, the other is for firewalled.
    • Upgrade to NDOS Timing System, doc-db 5871 ; master-slave configuration will be installed after this meeting.
  • Jon complained about dcm names and is concerned that we are not tracking when dcms go in and out of the detector.

DAQ/Electronics Status & Plans

  • Denis is working on determining if some TECCs can be recovered.
  • Peter: want to roll forward on message facility, but maybe not this week.
  • Jon: started making (simple) changes in RC for passthrough option
  • Andrew: working on making TDUControl a RCClient
  • Jon: started working on bringing local NDOS dB up and running

DCS Status & Plans

  • Athans:
    • question re: workshop on Friday: what are peoples' thoughts on how the meeting went?
      • Leon: seemed to go ok, got some commitments from FNAL personnel
      • Andrew: we will get ~100% of Ron's time for the next few weeks
      • Peter: why are we having Ron work on EPICS now?
      • Leon: this work serves 2 roles: resolves an immediate need, but also is a test of what we need in the future
    • will be meeting with Jim K.'s group this afternoon around 2:30 pm to discuss this further
    • still seeing problems with DDS suddenly stopping without any errors
    • also, dcmapplication seems to be blocking reading of registers when it is started up after Athan's program is running. Sue is looking into this.


  • None